# Advancing Video Anomaly Detection: A Concise Review and a New Dataset

### v3

@@ 1 Introduction
Surveys. Advances in VAD have also led to some comprehensive surveys contributing significantly to the literature [78, 1, 76, 17, 32]. A survey on deep learning models for VAD can be found in [78]. A comprehensive study in [1] highlights challenging issues in VAD, such as varying environments, the complexity of human activities, the ambiguous nature of anomalies, and the absence of appropriate datasets. A systematic review in [76] discusses various potential challenges, opportunities, and provides guidance for future research. A recent work [17] offers a comprehensive benchmark for understanding VAD causation using Large Language Models (LLMs) for text annotations on human-related anomalies. While we focus on static cameras, a survey on VAD in dynamic scenes captured by moving cameras can be found in [32]. Although these surveys are comprehensive, they are not portable and lightweight. Our lightweight review offers several benefits: (i) it provides a quick reference for researchers and practitioners, making it easier to get up to speed on VAD without sifting through extensive details; (ii) it enhances accessibility for a broader audience, including newcomers to the field, by allowing a quick grasp of essential concepts; and (iii) it offers focused guidance by distilling critical information into clear, actionable insights.
@@ 2.1 Dataset deficiencies and biases
From handcrafted to learned features. Early VAD methods use traditional techniques such as background subtraction, optical flow, and handcrafted feature extraction [50, 85, 72, 81], relying on appearance, motion, and texture to model motions and crowds [10, 111, 110, 5, 64]. However, these features are often insufficient due to the low resolution of benchmarks and limited training sources [3, 51, 86, 41], e.g., Subway, UMN, UCSD Ped and CUHK Avenue, etc. Subsequent works [33, 38, 58, 104, 116] explore the combination of local/global features, spatial/temporal normalcy, etc. Nevertheless, these methods are ineffective when applied to different camera views and cannot adapt to unseen anomalies, when ShanghaiTech (13 camera views) is introduced. Consequently, emerging methods predominantly investigate the use of learned features, eliminating the need for handcrafted features and making them more adaptable to various camera viewpoints and new scenarios [42]. These methods mainly focus on creating new architectures or modules tailored to specific problems, and they can be broadly classified into four categories: reconstruction-based [4, 25, 31], prediction-based [39, 42, 23, 7, 68, 88], using classifiers  [73, 53, 74, 75, 49, 31, 109], and scoring  [35, 80, 30, 19, 44], ranging from simple architectures [92] to complex unified approaches [13].
@@
Challenges of learned features. Most deep learning-based methods still carefully consider several aspects, such as visual appearance and motion, of human action [89, 39, 31, 98, 23, 93, 101, 91, 90, 12]. While these works highlight the importance of appearance and motion in detecting mostly human-related anomalies [23, 29, 31], non-human-related anomalies remain under-explored due to the lack of solid benchmarks, even when the UCF-Crime dataset (1 non-human-related anomaly) is introduced. Recent works [4] have delved into creating end-to-end deep learning approaches and unified architectures rather than using separate modules or components in a traditional pipeline, as end-to-end solutions are easily accessible, usable, and deployable. However, deep learning solutions require a significant amount of training data, posing a significant concern, especially considering older and smaller datasets such as UCSD Ped [86] and CUHK Avenue [41]. Although efforts have been made in collecting large-scale datasets [79, 107] such as UCF-crime (video- and frame-level annotations for training and testing, respectively), an important issue lies in laborious video annotation, which is one of the main reasons why there haven’t been as many large-scale VAD datasets published yet despite tons of data being publicly available on video sharing sites. A recent emerging few-shot learning framework [22] also encourages researchers to start looking at few-shot VAD models, due to (i) its fast adaptation to novel camera viewpoints/scenarios, (ii) relieving the training data hungry issue, and (iii) its huge potential in real-world applications. Since then, even smaller datasets have proven to be helpful [42, 45].
@@
Expanding modalities for human-related anomaly detection. Human-related anomaly detection solutions are not limited to the use of RGB videos. To efficiently extract motions, optical flow, precomputed on RGB videos, has been widely used as temporal or motion information for VAD. Due to the heavy computational cost of optical flow and the redundant nature of RGB videos, researchers have started to explore alternative data modalities for efficient feature extraction [103]. Human pose estimation frameworks, like OpenPose [8], have made human skeleton sequences readily available from RGB videos. This not only introduces a new data modality but also addresses privacy concerns in human-related anomaly detection. The lightweight nature of skeletons [60, 34, 96] and informative spatio-temporal sequences have fostered many skeletal anomaly detection solutions [52, 29]. For example, [52] uses 2D human skeleton motions to detect anomalous human behavior in surveillance footage. Recently, LLMs and pretrained video caption models provide rich descriptions and prompts for video contents to assist VAD tasks [17, 46, 115]. Since then, multi-modal datasets have begun to emerge, with representative examples such as CUVA (RGB videos with text descriptions) in 2024.
@@ 4.1 Generalizability and adaptability
Table 3: Evaluations on cross-scenario setups. We use FSAD [42] and SA2D (ours) for training on ShanghaiTech (ShT) and MSAD, respectively.
@@
MSAD UCSD Ped2 70.35 65.74
CUHK Avenue 79.57 84.49
@@
Setup. We first show the superior performance of our scenario-adaptive model by comparing SA2D with the FSAD model [42] on CUHK Avenue. To demonstrate the enhanced anomaly detection performance of our MSAD dataset, we train two adaptive models: one using ShanghaiTech, and the other leveraging our MSAD. For all experiments, we maintain $N\!=\!7$ and $K\!=\!10$ to ensure a fair comparison. Our training process spans over 1500 epochs. Throughout all evaluations, we consider both Micro and Macro AUC scores. We conduct two sets of experiments: (1) In the single-scenario / cross-view evaluation, the model is trained and tested on the same scenario. To facilitate a more comprehensive comparison with the view-adaptive anomaly detection model [42], we partition the ShanghaiTech dataset into 7 scenes (views), specifically, scene 2, 4, 7, 9, 10, 11, and 12 for training as in [42]. Subsequently, the remaining 5 camera views are individually used for testing purposes. [^3] There are no test videos for scene 13, so we only consider 5 views during testing. (2) In the cross-scenario evaluation, the model is trained on one scenario and tested on a completely different one. For model training, we use either ShanghaiTech or our MSAD.
@@
Cross-scenario evaluation. Our model, trained on MSAD, demonstrates superior performance compared to the view-adaptive model (see Table 3). For instance, on CUHK Avenue, our model outperforms the view-adaptive model by 9.6% and 6.2% for Micro and Macro evaluation metrics, respectively. Furthermore, our SA2D, trained and tested on MSAD, consistently achieves excellent performance. It’s important to note that our MSAD test set is distinct from the MSAD training set, covering a diverse range of novel scenarios. These results underscore the robustness of our model in cross-scenario evaluations. However, we also observe a performance drop on ShanghaiTech ($v6$). The decline in performance is attributed to the dataset deviating from real-life scenarios, as it categorizes biking and driving as anomalies, a classification that does not align with reality.
@@ Appendix A Dataset collection process
We have obtained permission from YouTube, Itemfix, and Bilibili to use their video data for non-commercial purposes. To facilitate the data collection, we use a script to automatically download relevant videos from these platforms and trim them into short clips. We also release these scripts along with our dataset for interested researchers. In our collection process, most videos are collected from YouTube (717 videos), and only a few videos are collected from BiliBili and ItemFix (3 videos). We ensure that all the videos are selected and used in accordance with the Fair Dealing provisions of the Australian Copyright Act 1968. Our approach is designed to adhere strictly to these legal frameworks, ensuring that the use of copyrighted material is only limited to research purposes. This careful consideration helps to protect the rights of the original content creators while allowing us to utilize the material in a legally compliant manner.
@@
Identifiable information. To eliminate any identifiable information, we consider the following options. (i) Face blurring: Since in surveillance video footage, the facial regions of individuals are often small and blurry, or the camera angles typically capture side or rear views, it is challenging to extract identifiable information about individuals. To eliminate any identifiable information that might be visible, we apply an automatic blurring script [^4] https://github.com/ORB-HD/deface to anonymize all the faces and car licenses in the videos of the MSAD dataset, then we check the whole dataset manually to ensure that all the identity information is almost eliminated. The blurred videos have limited impact to the detection performance while reducing the invasion of privacy. We provide the blurred version of our MSAD dataset. (ii) Extracted features: We have also released the extracted video features with different backbones, e.g., I3D, SwinTransformer, etc., to the research community to promote privacy preservation. We have noticed that some weakly-supervised methods in recent years use extracted video features as inputs [83, 14, 13, 108]. Training and evaluation with extracted video features is lightweight while protecting privacy. (iii) Deepfake technique: We also explore using deepfake techniques to remove identifiable features from the original videos, aiming to preserve characteristics such as age and gender while minimizing information loss compared to face blurring. However, we encounter challenges due to the diverse resolutions and camera angles, which make deepfake difficult. We plan to further investigate this approach in future work.
@@ Appendix B Further dataset details
Human-related Assault Assault on street Crime
@@
Non-human-related Explosion Street explosion Emergency
@@ Appendix E Further discussions
2023 VAD-CLIP [108] Language-visual model RGB+text - - - 88.0 84.5
@@
Enhancing anomaly understanding. Most existing methods primarily focus on detection accuracy [55], while only a few works address interpretable anomaly detection [17, 65]. Since anomalous behaviors are scene-dependent [6], it is essential for detection methods to understand the underlying rules and constraints of the scene in order to comprehend the causes and consequences of anomalies. Although some existing works attempt to integrate visual and semantic information to understand anomalies [13, 59, 108, 114], the guidance provided by semantic information for anomaly detection is still insufficient. Therefore, continuing to explore new methods, e.g., Large Language Models (LLMs) [17, 46, 115]), that enable models to better understand the semantic information of anomalies is crucial for advancing the field.
@@
With vision-language models. Exploring advanced deep learning and generative models to better capture the complexity of real-world scenarios is of vital importance. In particular, using pretrained large-scale vision-language models (VLMs) allows for a better understanding of anomalies in videos and produces corresponding descriptions. Notably, researchers have achieved state-of-the-art performance on multiple VAD tasks by using CLIP [61] to extract deep features, followed by a lightweight trainable network. Modeling both visual and language information for multimodal fusion, contextual understanding, handling complex scenarios, and providing explanatory descriptions would be an interesting future direction.
@@ Appendix F Further evaluations
Concerns on spurious correlations in anomalous event detection. We acknowledge the potential for a model to learn spurious correlations related to gender, race, or location when training on our dataset, especially in action classes like assault, fighting, and robbery, where faces and locations may be visible. We recognize the risk that a black-box video anomaly detection model could inadvertently learn such correlations, leading to biased or unfair outcomes. To address this, we propose the following recommendations and safeguards: (i) implementing additional anonymization techniques in the dataset, such as blurring or masking faces and identifiable features in the training data, with our blurred version of MSAD provided for researchers; (ii) using balanced sampling strategies and data augmentation techniques to ensure diverse representation across demographics and locations, minimizing the risk of unintended associations; (iii) incorporating post-training bias detection methods to identify learned spurious correlations, applying techniques such as fairness-aware learning [36] to mitigate these biases during training and evaluation; (iv) encouraging the use of explainable AI methods [16, 17] in anomaly detection to provide insights into the model’s decision features, promoting transparency to identify and address spurious correlations early; and (v) proposing continuous monitoring and evaluation of the model on diverse, representative test sets to ensure it generalizes well and avoids biased behavior in different contexts. In future, we plan to incorporate these recommendations more explicitly and explore further safeguards to ensure the ethical use of our dataset in training video anomaly detection models.
@@ Appendix I Potential societal impact
Our new dataset will enhance robust VAD, providing support and guidance to improve the safety and security of our communities.
However, since our dataset contains human identity information, it is crucial to explore methods that balance VAD performance while removing personally identifiable information, such as faces, poses, or gaits of individuals, even though we have obtained permissions from YouTube, Itemfix, and Bilibili to use their videos for research purposes. Our long-term plan is to explore new data modalities, such as the use of audio and skeleton sequences that do not contain human identity information, to advance VAD.

### v4

@@ 1 Introduction
Surveys. Advances in VAD have also led to some comprehensive surveys contributing significantly to the literature [18, 84, 60, 1, 81, 34, 115, 56]. A survey on deep learning models for VAD can be found in [84]. A comprehensive study in [1] highlights challenging issues in VAD, such as varying environments, the complexity of human activities, the ambiguous nature of anomalies, and the absence of appropriate datasets. A systematic review in [81] discusses various potential challenges, opportunities, and provides guidance for future research. A recent work [18] offers a comprehensive benchmark for understanding VAD causation using Large Language Models (LLMs) for text annotations on human-related anomalies. While we focus on static cameras, a survey on VAD in dynamic scenes captured by moving cameras can be found in [34]. Although these surveys are comprehensive, they are not portable and lightweight. Our lightweight review offers several benefits: (i) it provides a quick reference for researchers and practitioners, making it easier to get up to speed on VAD without sifting through extensive details; (ii) it enhances accessibility for a broader audience, including newcomers to the field, by allowing a quick grasp of essential concepts; and (iii) it offers focused guidance by distilling critical information into clear, actionable insights.
@@ 2.1 Dataset deficiencies and biases
From handcrafted to learned features. Early VAD methods use traditional techniques such as background subtraction, optical flow, and handcrafted feature extraction [55, 92, 82, 77, 88], relying on appearance, motion, and texture to model motions and crowds [10, 122, 121, 5, 69]. However, these features are often insufficient due to the low resolution of benchmarks and limited training sources [3, 54, 93, 45], e.g., Subway, UMN, UCSD Ped and CUHK Avenue, etc. Subsequent works [35, 41, 63, 111, 128] explore the combination of local/global features, spatial/temporal normalcy, etc. Nevertheless, these methods are ineffective when applied to different camera views and cannot adapt to unseen anomalies, when ShanghaiTech (13 camera views) is introduced. Consequently, emerging methods predominantly investigate the use of learned features, eliminating the need for handcrafted features and making them more adaptable to various camera viewpoints and new scenarios [46]. These methods mainly focus on creating new architectures or modules tailored to specific problems, and they can be broadly classified into four categories: reconstruction-based [4, 26, 33], prediction-based [43, 46, 24, 7, 73, 95], using classifiers  [78, 58, 79, 80, 53, 33, 117], and scoring  [37, 86, 32, 20, 48], ranging from simple architectures [99] to complex unified approaches [13].
@@
Challenges of learned features. Most deep learning-based methods still carefully consider several aspects, such as visual appearance and motion, of human action [96, 43, 33, 105, 24, 100, 108, 98, 15, 97, 12]. While these works highlight the importance of appearance and motion in detecting mostly human-related anomalies [24, 31, 33], non-human-related anomalies remain under-explored due to the lack of solid benchmarks, even when the UCF-Crime dataset (1 non-human-related anomaly) is introduced. Recent works [4] have delved into creating end-to-end deep learning approaches and unified architectures rather than using separate modules or components in a traditional pipeline, as end-to-end solutions are easily accessible, usable, and deployable. However, deep learning solutions require a significant amount of training data, posing a significant concern, especially considering older and smaller datasets such as UCSD Ped [93] and CUHK Avenue [45]. Although efforts have been made in collecting large-scale datasets [85, 114] such as UCF-crime (video- and frame-level annotations for training and testing, respectively), an important issue lies in laborious video annotation, which is one of the main reasons why there haven’t been as many large-scale VAD datasets published yet despite tons of data being publicly available on video sharing sites. A recent emerging few-shot learning framework [23] also encourages researchers to start looking at few-shot VAD models, due to (i) its fast adaptation to novel camera viewpoints/scenarios, (ii) relieving the training data hungry issue, and (iii) its huge potential in real-world applications. Since then, even smaller datasets have proven to be helpful [46, 49].
@@
Expanding modalities for human-related anomaly detection. Human-related anomaly detection solutions are not limited to the use of RGB videos. To efficiently extract motions, optical flow, precomputed on RGB videos, has been widely used as temporal or motion information for VAD [125]. Due to the heavy computational cost of optical flow and the redundant nature of RGB videos, researchers have started to explore alternative data modalities for efficient feature extraction [110]. Human pose estimation frameworks, like OpenPose [8], have made human skeleton sequences readily available from RGB videos. This not only introduces a new data modality but also addresses privacy concerns in human-related anomaly detection. The lightweight nature of skeletons [65, 36, 103] and informative spatio-temporal sequences have fostered many skeletal anomaly detection solutions [57, 31]. For example, [57] uses 2D human skeleton motions to detect anomalous human behavior in surveillance footage. Recently, LLMs and pretrained video caption models provide rich descriptions and prompts for video contents to assist VAD tasks [18, 50, 127]. Since then, multi-modal datasets have begun to emerge, with representative examples such as CUVA (RGB videos with text descriptions) in 2024.
@@ 4 Experiment
Table 3: Cross-scenario evaluations using FSAD [46] and SA2D (ours) trained on ShT and MSAD.
@@
MSAD Ped2 70.35 65.74
CUHK 79.57 84.49
@@
4.1 Generalizability and adaptability
Setup. We first show the superior performance of our scenario-adaptive model by comparing SA2D with the FSAD model [46] on CUHK Avenue. To demonstrate the enhanced anomaly detection performance of our MSAD dataset, we train two adaptive models: one using ShanghaiTech, and the other leveraging our MSAD. For all experiments, we maintain $N\!=\!7$ and $K\!=\!10$ to ensure a fair comparison. Our training process spans over 1500 epochs. Throughout all evaluations, we consider both Micro and Macro AUC scores. We conduct two sets of experiments: (1) In the single-scenario / cross-view evaluation, the model is trained and tested on the same scenario. To facilitate a comparison with the view-adaptive anomaly detection model [46], we partition the ShanghaiTech dataset into 7 scenes (views), specifically, scene 2, 4, 7, 9, 10, 11, and 12 for training as in [46]. Subsequently, the remaining 5 camera views are individually used for testing purposes. [^3] There are no test videos for scene 13, so we only consider 5 views during testing. (2) In the cross-scenario evaluation, the model is trained on one scenario and tested on a completely different one. For model training, we use either ShanghaiTech or our MSAD.
@@ 4.1 Generalizability and adaptability
Cross-scenario evaluation. Our model, trained on MSAD, demonstrates superior performance compared to the view-adaptive model (see Table 3, Ped2 and CUHK refer to UCSD Ped2 and CUHK Avenue datasets). For instance, on CUHK Avenue, our model outperforms the view-adaptive model by 9.6% and 6.2% for Micro and Macro evaluation metrics, respectively. Furthermore, our SA2D, trained and tested on MSAD, consistently achieves excellent performance. It’s important to note that our MSAD test set is distinct from the MSAD training set, covering a diverse range of novel scenarios. These results underscore the robustness of our model in cross-scenario evaluations. However, we also observe a performance drop on ShanghaiTech ($v6$). The decline in performance is attributed to the dataset deviating from real-life scenarios, as it categorizes biking and driving as anomalies, a classification that does not align with reality.
@@ Appendix A Dataset collection process
To facilitate the data collection, we use a script to automatically download relevant videos from YouTube, Itemfix, and Bilibili and trim them into short clips. We ensure that all the videos are selected and used in accordance with the fair dealing provisions of the Australian Copyright Act 1968. Our approach is designed to adhere strictly to these legal frameworks, ensuring that the use of copyrighted material is only limited to research purposes. This careful consideration helps to protect the rights of the original content creators while allowing us to utilize the material in a legally compliant manner.
@@
Identifiable information. To eliminate any identifiable information, we consider the following options. (i) Face blurring: Since in surveillance video footage, the facial regions of individuals are often small and blurry, or the camera angles typically capture side or rear views, it is challenging to extract identifiable information about individuals. To eliminate any identifiable information that might be visible, we apply an automatic blurring script [^4] https://github.com/ORB-HD/deface to anonymize all the faces and car licenses in the videos of the MSAD dataset, then we check the whole dataset manually to ensure that all the identity information is almost eliminated. The blurred videos have limited impact to the detection performance while reducing the invasion of privacy. We provide the blurred version of our MSAD dataset. (ii) Extracted features: We have also released the extracted video features with different backbones, e.g., I3D, SwinTransformer, etc., to the research community to promote privacy preservation. We have noticed that some weakly-supervised methods in recent years use extracted video features as inputs [39, 90, 14, 13, 116]. Training and evaluation with extracted video features is lightweight while protecting privacy. (iii) Deepfake technique: We also explore using deepfake techniques to remove identifiable features from the original videos, aiming to preserve characteristics such as age and gender while minimizing information loss compared to face blurring. However, we encounter challenges due to the diverse resolutions and camera angles, which make deepfake difficult. We plan to further investigate this approach in future work.
@@ Appendix B Further dataset details
Human-related anomaly Assault Assault on street Crime
@@
Non-human-related anomaly Explosion Street explosion Emergency
@@ Appendix E Further discussions
2023 HSC [87] Semantic contrastive learning RGB 98.1 93.7 83.4 - -
2023 USTN-DSC [120] Video event restoration RGB 98.1 89.9 73.8 - -
@@
2023 VAD-CLIP [116] Vision-language model RGB+text - - - 88.0 84.5
@@
Enhancing anomaly understanding. Most existing methods primarily focus on detection accuracy [60], while only a few works address interpretable anomaly detection [18, 70]. Since anomalous behaviors are scene-dependent [6], it is essential for detection methods to understand the underlying rules and constraints of the scene in order to comprehend the causes and consequences of anomalies. Although some existing works attempt to integrate visual and semantic information to understand anomalies [13, 64, 116, 126], the guidance provided by semantic information for anomaly detection is still insufficient. Therefore, continuing to explore new methods, e.g., Large Language Models (LLMs) [18, 50, 127], that enable models to better understand the semantic information of anomalies is crucial for advancing the field.
@@
With vision-language models. Exploring advanced deep learning and generative models to better capture the complexity of real-world scenarios is of vital importance. In particular, using pretrained large-scale vision-language models (VLMs) enhances the understanding of anomalies in videos and facilitates a more coherent reasoning process [119, 28, 50]. Notably, researchers have achieved state-of-the-art performance on multiple VAD tasks by using CLIP [66] to extract deep features, followed by a lightweight trainable network [116]. Modeling both visual and language information for multimodal fusion, contextual understanding, handling complex scenarios, and providing explanatory descriptions would be an interesting future direction.
@@ Appendix F Further evaluations
Concerns on spurious correlations in anomalous event detection. We acknowledge the potential for a model to learn spurious correlations related to gender, race, or location when training on our dataset, especially in action classes like assault, fighting, and robbery, where faces and locations may be visible. We recognize the risk that a black-box video anomaly detection model could inadvertently learn such correlations, leading to biased or unfair outcomes.
To address this, we propose the following recommendations and safeguards: (i) implementing additional anonymization techniques in the dataset, such as blurring or masking faces and identifiable features in the training data, with our blurred version of MSAD provided for researchers; (ii) using balanced sampling strategies and data augmentation techniques to ensure diverse representation across demographics and locations, minimizing the risk of unintended associations; (iii) incorporating post-training bias detection methods to identify learned spurious correlations, applying techniques such as fairness-aware learning [38] to mitigate these biases during training and evaluation; (iv) encouraging the use of explainable AI methods [17, 18] in anomaly detection to provide insights into the model’s decision features, promoting transparency to identify and address spurious correlations early; and (v) proposing continuous monitoring and evaluation of the model on diverse, representative test sets to ensure it generalizes well and avoids biased behavior in different contexts.
In future, we plan to incorporate these recommendations more explicitly and explore further safeguards to ensure the ethical use of our dataset in training video anomaly detection models.
@@ Appendix I Potential societal impact
Our MSAD dataset is designed to advance video anomaly detection and offers crucial support for enhancing community safety and security in surveillance applications.
To protect privacy, we have applied anonymization techniques (see Sec. A) to remove identifiable human information, such as faces, poses, and gaits. However, privacy remains a critical area for ongoing research, especially in sensitive environments like homes, hospitals, and care facilities, where it is essential to balance effective detection with robust privacy safeguards.
In the future, we aim to explore new data modalities, including audio and skeletal sequences that inherently lack identifiable information, to further improve video anomaly detection.